OpenStudy (anonymous):

Solve 5x2 = -30x - 65

OpenStudy (muzzack):

Solution Steps 1. Multiply Terms 5x2=-30x−65 2. Move Variables to Left 5x2+30x=-65 3. Apply Quadratic Formula x=−30±302−4*5*65√2*5 Forked Equation 1 x=−30+302−4*5*65√2*5 Forked Equation 2 x=−30−302−4*5*65√2*5 Now Solving Equation 1 x=−30+302−4*5*65√2*5 4. Negate x=-30+302−4*5*65√2*5 5. Simplify Exponents x=-30+900−4*5*65√2*5 6. Multiply Terms x=-30+900−1300√10 7. Subtract Constants x=-30+-400√10 The equation has an imaginary solution! End of Equation 1 Now Solving Equation 2 x=−30−302−4*5*65√2*5 8. Negate x=-30−302−4*5*65√2*5 9. Simplify Exponents x=-30−900−4*5*65√2*5 10. Multiply Terms x=-30−900−1300√10 11. Subtract Constants x=-30−-400√10 The equation has an imaginary solution!

OpenStudy (anonymous):

all the answers have an i on the end of them?

OpenStudy (muzzack):

lol idk

OpenStudy (anonymous):

x = -3 ± 2i x = -3 ± 4i x = -6 ± 2i x = -6 ± 4i

OpenStudy (anonymous):

@Muzzack